Supplemental Operations Report – November 19th

Name of person filing report: Shannon Rupert
Reason for Report: Routine
Non-nominal systems: Both septic tanks were pumped out on Nov 11th.
Action taken for non-nominal systems: N/A
Generator: Oil, oil filter and air filter changed Nov 12th. We are running it about 13 hours a day right now, which is normal for this time of year.
ScienceDome Dual Split: off
Solar— Nominal.
Solar— SOC Last 24 hours:
Max 100%
Min 68%
Avg 76.1%
Propane Reading, station tank – full, refilled on Nov 17, 2022
Propane Reading, director tank— full, refilled on Nov 17th
Propane Reading, intern tank— full, refilled on Nov 17th
Propane Reading, generator— full, refilled on Nov 17th
Ethanol Free Gasoline – 0 gallons
Water (static tank) – 250 gallons
Water in GreenHab – 150 gallons
Water in ScienceDome: 0 gallons
Water (Outpost tank) – 300 gallons
Hab toilet tank emptied: check regular operations reports
Sojourner rover used: yes
Hours: not noted
Beginning charge: 100
Ending charge: 100
Currently charging: yes
Notes on rovers: All rovers have been serviced this week. Battery contacts were cleaned, batteries were filled with distilled water, seats were patched. Opportunity has died in the field twice this season, once with each crew.
ATV’s Used: (Honda, 350.1, 350.2, 300): Honda
Reason for use: charging battery
Oil Added? no
ATV Fuel Used: 0 Gals
# Hours the ATVs were Used today: 0
Notes on ATVs: Currently waiting to be picked up for serving
HabCar used and why, where? Yes, to town
CrewCar used and why, where? Yes, to town and to Grand Junction
Luna used and why, where? No
General notes and comments: Nothing to report
Summary of internet: All accounts are nominal but I keep getting kicked off the network, which I find super annoying.
Summary of suits and radios: We replaced 6 of the new EVA radios with longer whip antennas but did not see a noticeable increase in clarity of communication. Next step is to get the addition to the repeater installed and see if that changes things.
Campus wide inspection, if action taken, what and why? Nothing to report
Summary of general operations: Nominal
Summary of Hab operations: Finished the winterizing of the pipes between static tank and Hab. Next fall we will replace this with a direct line from the tank to the building. But for this window it is heat taped and foam wrapped, so should be ok. The tank heater will be installed today as well.
Summary of Outpost operations: I carpeted my living/kitchen areas with leftover carpet squares. I’m fancy now! Sergii installed LED lights in the living area of his trailer for better working light. We purchased new lights for the 110V lights in each trailer but they won’t work in mine. We have an additional set we will install in the Hab shower room and toilet next weekend. Sergii took a CO monitor from the Hab because he did not have one in his trailer. The propane alarm is also not working in that trailer so we will be replacing that as well. We have used the new washing machine and are quite pleased with it. Sergii also added some holes in the floor under the sink to keep the kitchen pipes from freezing at night. He insulated some of the windows with silver bubble wrap insulation as well. It’s been colder than normal for this time of year.
Summary of GreenHab operations: Thermostat was replaced and is working correctly for the first time in a couple of seasons.
Summary of ScienceDome operations: Nothing to report
Summary of RAM operations: Nothing to report
Summary of any observatory issues: Nice to see the dome open most nights!
Summary of health and safety issues: Nothing to report
Questions, concerns, supplies needed and requests: n/a

Copyright © The Mars Society. All rights reserved. | Main Site